# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ E.4 Results of negative control experiments
We provide additional results for our varied-prefix-length negative control experiments (Section 5.2) on both Phi 4 (Appendix E.4.1) and non-training data that post-dates training cutoffs (Appendix E.4.2). While these experiments provide strong evidence for the validity of our procedure, we would want to run much more extensive experiments to responsibly lower $\tau_{\text{min}}$ or to make it model-specific. We defer this to future work, and here opt to pick a very conservative global $\tau_{\text{min}}{=}0.1\%$, which we discuss also in Section 5. The experiments that we run here are of a more substantial scale than the original probabilistic discoverable extraction paper by Hayes et al. [60]. That work performed one experiment on non-training data on a small Pythia model. The paper uses the results of this experiment as a basis to both validate the experiments on known training data in Pythia models, and to apply the methodology to much larger (and higher quality) Llama 1 and OPT models—i.e., to make claims about extraction of not-known-with-certainty training data, just as we do here with DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B. Hayes et al. [60] finds that “…the number of queries $n$ needed to generate unseen test data is orders of magnitude larger than for generating training data. [They] find that it is generally challenging to generate test data …This supports that, in [their] measurements of [probabilistic] discoverable extraction, matches between training-example targets and generated suffixes are almost surely due to memorization.” We observe the same pattern, though instead of discussing it with respect to the number of queries $n$, as elsewhere, we frame our discussion in terms of $p_{\bm{z}}$. In our results, the $p_{\bm{z}}$ that we observe on non-training data are many orders of magnitude lower than what we deem extraction success in our experiments on known training data. That is, our measurements for extraction are not erroneously capturing chance generation of non-training data. As in Hayes et al. [60], this gives us assurance that we can use our approach on models for which we do not know exactly what was included in the training data. We discuss this further in Appendices E.5 and E.6.
@@ E.4.1 Results on Phi 4
We run a large number of negative control experiments using Phi 4 [2]. As noted in Appendix D.1, this model was not trained on whole copyrighted books. It was trained on a mix of synthetic data, licensed websites, and public domain books. When running our sliding-window extraction procedure on this model for copyrighted books, we can expect (generally speaking, with important exceptions) that unique text from copyrighted books will not register as extractable with our measurements. For our main results using the sliding-window procedure with $100$-token sequences (Appendix D.3), we include Phi 4 as one of the $10$ models that we evaluate. Here, in Figure 118, we show additional results for experiments on varied prefixes and $4$ in-copyright books from Books3. We are able to extract non-trivial amounts of public domain books from this model (e.g., Appendix D.3.17). In contrast, we are not able to extract much unique text from in-copyright text. As elsewhere, we extract boilerplate text like copyright notices and publishers’ addresses. And we are also able to extract famous or otherwise popular quotes—either quoted in in-copyright books, or from the books themselves.We (painstakingly) verify that every sequence that we extract from Phi 4 can be easily found in multiple sources on the Internet. That is, the quotes from books that we extract from Phi 4 are either in the public domain, quoted widely on blogs or platforms like Goodreads, or both. This is indicative (though not definitive) that these quotes could have found their way into Phi 4’s training data from these online sources, rather from the books themselves. For further support, we can also compare the text at the book locations for which we extract these sequences to heatmaps other models. We note that, almost always, these sequences are extractable (sometimes with enormous probability) from all models we test. Again, while this is not definitive, it is suggestive that this text is highly duplicated in sources that are a part of common pre-training corpora—that they are not necessarily memorized because they are in Books3, but because of their duplicated inclusion in the training data from other sources. We also observe similar challenges—to a somewhat lesser extent—for our negative control experiments on books published after LLM training-date cutoffs. These books sometimes cite earlier sources—other books, news articles, social media posts, etc.—that pre-date the training-date cutoffs and are widely duplicated online. So, even this seemingly “clean” case also requires significant manual verification to confirm validity, which we discuss in Appendix E.4.2. Altogether, it is reasonable to conclude that these are true instances of extraction. These results reinforce that our procedure detects true memorization when it exists and does not spuriously flag text where memorization is impossible. Even so, these results show just how challenging it is to run completely clean negative controls for models trained on web-scraped data [75], for which duplicates of the same piece of text across various sources complicate what training-data membership means. A given book may not have been a member of the training data, but this does not mean that all text from that book was not. In these cases, it seems extraordinary likely that these sequences are members of the training data (and memorized). We discuss this further in Appendix E.6. With the exception of the copyright notice, we are unable to extract any sequences from This Is How You Lose Her [40]. The sequences that we extract from Beloved are all posted (multiple times) on platforms like Goodreads. We discuss some of these quotes in Section 6.2 [90, 91, 92]. Harry Potter and the Sorcerer’s Stone is excerpted and quoted in an extraordinary number of places on the Internet. We were able to find every extracted sequence on $3$ different web pages (after which we stopped counting).
@@
Figure 118: Phi 4 Negative Control: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [89], Harry Potter and the Sorcerer’s Stone [106], This Is How You Lose Her [40], and We Were Eight Years in Power [23]. Phi 4 was deliberately not trained on whole copyrighted books. We run probabilistic extraction on Llama 3.1 70B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens. We manually verified that every extractable sequence comes from highly duplicated text that can easily be found on the Internet; such sources reasonably were a part of Phi 4’s training data.
@@
We discuss When We Were Eight Years in Power [23] in Section 5.2. As noted there, regardless of prefix length, there are only $4$ areas—other than the copyright notice at the beginning of the book—that show extractable sequences. The sequences at the end of the book contain a list of ordered (page) numbers. This kind of list is a very common artifact from converting an epub file to txt—as was done for the collection of Books3 [104]. Like the copyright notice, such ordered lists of numbers are text that, while high probability, we can discount as being particularly interesting with respect to our negative controls. The other $3$ are all text that is not unique to book, but rather highly duplicated text from the Internet (some of which is in the public domain):
@@
  • $\sim\!176$K: An exact quote from Malcom X’s speech, “Who Taught You To Hate Yourself?” (May 5, 1962) [137]
@@
  • $\sim\!280$K: An exact quote of Deuteronomy 15:125 from the Bible – King James Version (KJV) [124].
@@
  • $\sim\!546$K: A exact quote from Barack Obama’s 2004 DNC Keynote: “hope of a young naval lieutenant bravely patrolling the Mekong Delta; the hope of a mill worker’s son who dares to defy the odds; the hope of a skinny kid with a funny name who believes that America has a place for” [95]
@@
Figure 119: Post-Training Cutoff Negative Control: We measure the probability of generating target $50$-token suffixes for $50$-token prefixes (using top-$k$, as elsewhere) for $4$ books published in 2025: Careless People [136], Great Big Beautiful Life [63], The Emperor of Gladness [130], and The Society of Unknowable Objects [12]. All of these books were published after the release dates for all models we evaluate. These experiments serve as a validity check that our measurements for probabilistic extraction are capturing true instances of extraction. By definition, no unique sequences in these books can be memorized because they are not members of the training data; we should not be able to (and do not) obtain high, non-zero probabilities—probabilities that we would consider to be successful instances of extraction for known or suspected training data. Note that, in contrast to our other heatmaps, we set a more permissive (i.e., lower) minimum for $p_{\bm{z}}$ ($p_{\bm{z}}\geq 0.00000001\%$, or $1$ in $10$ billion generations), in order to see if there are sequences that can be generated even with fairly low probability that we would not count as successful extraction.
@@ E.4.2 Results on books from after training cutoffs
We run additional negative control experiments on $4$ books that post-date the training of all models. As a result, these books could not have been included in the training data; we similarly should not expect to extract unique text from them for the different models we evaluate. However, as with our experiments on Phi 4, there are again cases where boilerplate text and popular quotes from other sources come up in our measurements. Nevertheless, we are not able to extract any text unique to any of these $4$ books for any model we test. In Figure 119, we show results for all $4$ books using $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes) for the $10$ models that we include for our sliding-window experiments on Books3 books (Appendix D). Note that in Figure 119, we expand the range of $p_{\bm{z}}$ that we plot to orders of magnitude below $\tau_{\text{min}}{=}0.1\%$. Even so, the only regions that register with non-zero probability are a select few that surpass $\tau_{\text{min}}$, and they are duplicated generic text—in this case, the copyright notice—and some references drawn from pre-2025 sources. For example, for The Emperor of Gladness [130], we extract a verbatim quote from the Bible (Mark 5:13). For Great Big Beautiful Life [63] and The Society of Unknowable Objects [12], we only extract boilerplate text. For Careless People [136], we extract the “careless people” quote from The Great Gatsby [46] (Figure 1). It is quoted at the beginning of the book (e.g., $p_{\bm{z}}{=}2.8\%$ for Llama 3.1 70B):
@@
This is what we register as extracted close to the 0-character position in the heatmaps. (We highlight the portion from The Great Gatsby in blue.) The book also quotes Mark Zuckerberg in several locations—quotes that we confirm are often drawn from Facebook posts that are widely duplicated on social media platforms and in online newspaper articles. These make up the remaining extracted sequences. For example, one quote is from a 2016 post about a SpaceX rocket explosion [133] (e.g., $p_{\bm{z}}{=}13.5\%$ for Llama 1 65B):
@@
  • need to reject—a group of extremists trying to silence the voices and opinions of everyone else around the world. I won’t let that happen on Facebook. I’m committed to building a service where you can speak freely without fear of violence [88]
Mark Zuckerberg’s manifesto is also quoted [116]. We know the The Great Gatsby quote is in Books3, and widely quoted elsewhere. Similar to our observations for Phi 4, it is very reasonable that these other highly duplicated quotes are similarly in the training data, and that these are true instances of extraction.Overall, consistent with Hayes et al. [60], we do not observe chance generation of non-training data until we significantly lower the probability threshold. In Figure 119, we observe no such instances for $p_{\bm{z}}\geq 10^{-8}\%$, which is $7$ orders of magnitude smaller than our chosen $\tau_{\text{min}}$. (This is an upper bound; this choice was arbitrary for this plot.) We also include the last book, Great Big Beautiful Life [63], in our varied prefix experiments as another point of comparison. We discuss these results for Llama 3.1 70B in Section 5.2, and provide the full set of results for $5$ models in Figure 120. We extract no sequences (beyond boilerplate text) from any of the $5$ models we test. In Figure 7, we show an extended $p_{\bm{z}}$ range to highlight how low we need to go beneath $\tau_{\text{min}}$ to register chance generation of non-training data. We extend to $10^{-14}\%$$13$ orders of magnitude below $\tau_{\text{min}}{=}0.1\%$—and find exactly $1$ sequence that surpasses this threshold for 200200800800-token prefixes. For example, for $800$-token prefixes, we register $1$ sequence with $p_{\bm{z}}{=}6.06\times 10^{-14}\%$. We report this as on the order of $10^{-14}\%$ in the main text because that is all we can reliably do for a number this small. The mantissa ($6.06$) is unreliable at this magnitude. This is because, even though we perform all $\log$ probability computations on CPU in full float64 precision, producing the logits during inference on GPU (that we post-process on CPU) involves mixed-precision computations with the model weights/activations in bfloat16 [53] and accumulation operations (matrix multiplications, normalization) in float32 [64]. Each computation introduces small rounding errors that accumulate and propagate to the final logits. When we exponentiate to to get probabilities, these errors result in only the exponent (not the mantissa) being reliable.
@@
Figure 120: Post-Training Cutoff Negative Control: Sliding-window procedure for various prefix lengths. We show results for Great Big Beautiful Life [63], which was published in spring 2025—after the release of all of the models we test. (The newest model is from the Qwen 2.5 series, which was released in late September 2024.) This book is not in any of the models’ training data. We run probabilistic extraction on (a) Llama 3.1 70B, (b) Llama 3.1 8B, (c) Llama 2 13B, (d) Phi 4, and (e) Qwen 2.5 72B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. We are unable to extract any unique (i.e., non-copyright notice) sequences using any model or prefix length.

### v4

@@ E.4 Results of negative control experiments
We provide additional results for our varied-prefix-length negative control experiments (Section 5.2) on both Phi 4 (Appendix E.4.1) and non-training data that post-dates training cutoffs (Appendix E.4.2). While these experiments provide strong evidence for the validity of our procedure, we would want to run much more extensive experiments to responsibly lower $\tau_{\text{min}}$ or to make it model-specific. We defer this to future work, and here opt to pick a very conservative global $\tau_{\text{min}}{=}0.1\%$, which we discuss also in Section 5. The experiments that we run here are of a more substantial scale than the original probabilistic discoverable extraction paper by hayes2025measuringmemorizationlanguagemodels. That work performed one experiment on non-training data on a small Pythia model. The paper uses the results of this experiment as a basis to both validate the experiments on known training data in Pythia models, and to apply the methodology to much larger (and higher quality) Llama 1 and OPT models—i.e., to make claims about extraction of not-known-with-certainty training data, just as we do here with DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B.
hayes2025measuringmemorizationlanguagemodels finds that “…the number of queries $n$ needed to generate unseen test data is orders of magnitude larger than for generating training data. [They] find that it is generally challenging to generate test data …This supports that, in [their] measurements of [probabilistic] discoverable extraction, matches between training-example targets and generated suffixes are almost surely due to memorization.” We observe the same pattern, though instead of discussing it with respect to the number of queries $n$, as elsewhere, we frame our discussion in terms of $p_{\bm{z}}$. In our results, the $p_{\bm{z}}$ that we observe on non-training data are many orders of magnitude lower than what we deem extraction success in our experiments on known training data. That is, our measurements for extraction are not erroneously capturing chance generation of non-training data. As in hayes2025measuringmemorizationlanguagemodels, this gives us assurance that we can use our approach on models for which we do not know exactly what was included in the training data. We discuss this further in Appendices E.5 and E.6.
@@ E.4.1 Results on Phi 4
We run a large number of negative control experiments using Phi 4 [phi4]. As noted in Appendix D.1, this model was not trained on whole copyrighted books. It was trained on a mix of synthetic data, licensed websites, and public domain books. When running our sliding-window extraction procedure on this model for copyrighted books, we can expect (generally speaking, with important exceptions) that unique text from copyrighted books will not register as extractable with our measurements.
For our main results using the sliding-window procedure with $100$-token sequences (Appendix D.3), we include Phi 4 as one of the $10$ models that we evaluate. Here, in Figure 118, we show additional results for experiments on varied prefixes and $4$ in-copyright books from Books3. We are able to extract non-trivial amounts of public domain books from this model (e.g., Appendix D.3.17). In contrast, we are not able to extract much unique text from in-copyright text. As elsewhere, we extract boilerplate text like copyright notices and publishers’ addresses. And we are also able to extract famous or otherwise popular quotes—either quoted in in-copyright books, or from the books themselves.
We (painstakingly) verify that every sequence that we extract from Phi 4 can be easily found in multiple sources on the Internet. That is, the quotes from books that we extract from Phi 4 are either in the public domain, quoted widely on blogs or platforms like Goodreads, or both. This is indicative (though not definitive) that these quotes could have found their way into Phi 4’s training data from these online sources, rather from the books themselves.
For further support, we can also compare the text at the book locations for which we extract these sequences to heatmaps other models. We note that, almost always, these sequences are extractable (sometimes with enormous probability) from all models we test. Again, while this is not definitive, it is suggestive that this text is highly duplicated in sources that are a part of common pre-training corpora—that they are not necessarily memorized because they are in Books3, but because of their duplicated inclusion in the training data from other sources.
We also observe similar challenges—to a somewhat lesser extent—for our negative control experiments on books published after LLM training-date cutoffs. These books sometimes cite earlier sources—other books, news articles, social media posts, etc.—that pre-date the training-date cutoffs and are widely duplicated online. So, even this seemingly “clean” case also requires significant manual verification to confirm validity, which we discuss in Appendix E.4.2.
Altogether, it is reasonable to conclude that these are true instances of extraction. These results reinforce that our procedure detects true memorization when it exists and does not spuriously flag text where memorization is impossible. Even so, these results show just how challenging it is to run completely clean negative controls for models trained on web-scraped data [lee2023explainers], for which duplicates of the same piece of text across various sources complicate what training-data membership means. A given book may not have been a member of the training data, but this does not mean that all text from that book was not. In these cases, it seems extraordinary likely that these sequences are members of the training data (and memorized). We discuss this further in Appendix E.6.
With the exception of the copyright notice, we are unable to extract any sequences from This Is How You Lose Her [loseher]. The sequences that we extract from Beloved are all posted (multiple times) on platforms like Goodreads. We discuss some of these quotes in Section 6.2 [belovedquote1, belovedquote2, belovedquote3]. Harry Potter and the Sorcerer’s Stone is excerpted and quoted in an extraordinary number of places on the Internet. We were able to find every extracted sequence on $3$ different web pages (after which we stopped counting).
@@
Figure 118: Phi 4 Negative Control: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. Phi 4 was deliberately not trained on whole copyrighted books. We run probabilistic extraction on Llama 3.1 70B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens. We manually verified that every extractable sequence comes from highly duplicated text that can easily be found on the Internet; such sources reasonably were a part of Phi 4’s training data.
@@
We discuss When We Were Eight Years in Power [wewere] in Section 5.2. As noted there, regardless of prefix length, there are only $4$ areas—other than the copyright notice at the beginning of the book—that show extractable sequences. The sequences at the end of the book contain a list of ordered (page) numbers. This kind of list is a very common artifact from converting an epub file to txt—as was done for the collection of Books3 [reisnerbooks3]. Like the copyright notice, such ordered lists of numbers are text that, while high probability, we can discount as being particularly interesting with respect to our negative controls. The other $3$ are all text that is not unique to book, but rather highly duplicated text from the Internet (some of which is in the public domain):
@@
  • $\sim\!176$K: An exact quote from Malcom X’s speech, “Who Taught You To Hate Yourself?” (May 5, 1962) [malcolmx]
@@
  • $\sim\!280$K: An exact quote of Deuteronomy 15:125 from the Bible – King James Version (KJV) [deuteronomy].
@@
  • $\sim\!546$K: A exact quote from Barack Obama’s 2004 DNC Keynote: “hope of a young naval lieutenant bravely patrolling the Mekong Delta; the hope of a mill worker’s son who dares to defy the odds; the hope of a skinny kid with a funny name who believes that America has a place for” [obama2004]
@@
Figure 119: Post-Training Cutoff Negative Control: We measure the probability of generating target $50$-token suffixes for $50$-token prefixes (using top-$k$, as elsewhere) for $4$ books published in 2025: Careless People [careless], Great Big Beautiful Life [greatbig], The Emperor of Gladness [emperorgladness], and The Society of Unknowable Objects [unknowableobjects]. All of these books were published after the release dates for all models we evaluate. These experiments serve as a validity check that our measurements for probabilistic extraction are capturing true instances of extraction. By definition, no unique sequences in these books can be memorized because they are not members of the training data; we should not be able to (and do not) obtain high, non-zero probabilities—probabilities that we would consider to be successful instances of extraction for known or suspected training data. Note that, in contrast to our other heatmaps, we set a more permissive (i.e., lower) minimum for $p_{\bm{z}}$ ($p_{\bm{z}}\geq 0.00000001\%$, or $1$ in $10$ billion generations), in order to see if there are sequences that can be generated even with fairly low probability that we would not count as successful extraction.
@@ E.4.2 Results on books from after training cutoffs
We run additional negative control experiments on $4$ books that post-date the training of all models. As a result, these books could not have been included in the training data; we similarly should not expect to extract unique text from them for the different models we evaluate. However, as with our experiments on Phi 4, there are again cases where boilerplate text and popular quotes from other sources come up in our measurements.
Nevertheless, we are not able to extract any text unique to any of these $4$ books for any model we test. In Figure 119, we show results for all $4$ books using $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes) for the $10$ models that we include for our sliding-window experiments on Books3 books (Appendix D). Note that in Figure 119, we expand the range of $p_{\bm{z}}$ that we plot to orders of magnitude below $\tau_{\text{min}}{=}0.1\%$. Even so, the only regions that register with non-zero probability are a select few that surpass $\tau_{\text{min}}$, and they are duplicated generic text—in this case, the copyright notice—and some references drawn from pre-2025 sources. For example, for The Emperor of Gladness [emperorgladness], we extract a verbatim quote from the Bible (Mark 5:13). For Great Big Beautiful Life [greatbig] and The Society of Unknowable Objects [unknowableobjects], we only extract boilerplate text.
For Careless People [careless], we extract the “careless people” quote from The Great Gatsby [gatsby] (Figure 1). It is quoted at the beginning of the book (e.g., $p_{\bm{z}}{=}2.8\%$ for Llama 3.1 70B):
@@
This is what we register as extracted close to the 0-character position in the heatmaps. (We highlight the portion from The Great Gatsby in blue.) The book also quotes Mark Zuckerberg in several locations—quotes that we confirm are often drawn from Facebook posts that are widely duplicated on social media platforms and in online newspaper articles. These make up the remaining extracted sequences. For example, one quote is from a 2016 post about a SpaceX rocket explosion [spacex] (e.g., $p_{\bm{z}}{=}13.5\%$ for Llama 1 65B):
[…]
    Suffix: hear that SpaceX’s launch failure destroyed our satellite that would have provided connectivity to so many entrepreneurs and everyone else across the continent. Fortunately, we have developed other technologies like Aquila that will connect people as well.
@@
  • need to reject—a group of extremists trying to silence the voices and opinions of everyone else around the world. I won’t let that happen on Facebook. I’m committed to building a service where you can speak freely without fear of violence [hebdo]
Mark Zuckerberg’s manifesto is also quoted [manifesto].
We know the The Great Gatsby quote is in Books3, and widely quoted elsewhere. Similar to our observations for Phi 4, it is very reasonable that these other highly duplicated quotes are similarly in the training data, and that these are true instances of extraction.
Overall, consistent with hayes2025measuringmemorizationlanguagemodels, we do not observe chance generation of non-training data until we significantly lower the probability threshold. In Figure 119, we observe no such instances for $p_{\bm{z}}\geq 10^{-8}\%$, which is $7$ orders of magnitude smaller than our chosen $\tau_{\text{min}}$. (This is an upper bound; this choice was arbitrary for this plot.)
We also include the last book, Great Big Beautiful Life [greatbig], in our varied prefix experiments as another point of comparison. We discuss these results for Llama 3.1 70B in Section 5.2, and provide the full set of results for $5$ models in Figure 120. We extract no sequences (beyond boilerplate text) from any of the $5$ models we test. In Figure 7, we show an extended $p_{\bm{z}}$ range to highlight how low we need to go beneath $\tau_{\text{min}}$ to register chance generation of non-training data. We extend to $10^{-14}\%$$13$ orders of magnitude below $\tau_{\text{min}}{=}0.1\%$—and find exactly $1$ sequence that surpasses this threshold for 200200800800-token prefixes.
For example, for $800$-token prefixes, we register $1$ sequence with $p_{\bm{z}}{=}6.06\times 10^{-14}\%$. We report this as on the order of $10^{-14}\%$ in the main text because that is all we can reliably do for a number this small. The mantissa ($6.06$) is unreliable at this magnitude. This is because, even though we perform all $\log$ probability computations on CPU in full float64 precision, producing the logits during inference on GPU (that we post-process on CPU) involves mixed-precision computations with the model weights/activations in bfloat16 [gdm_bf16_range] and accumulation operations (matrix multiplications, normalization) in float32 [hf_bf16_accum]. Each computation introduces small rounding errors that accumulate and propagate to the final logits. When we exponentiate to to get probabilities, these errors result in only the exponent (not the mantissa) being reliable.
@@
Figure 120: Post-Training Cutoff Negative Control: Sliding-window procedure for various prefix lengths. We show results for Great Big Beautiful Life [greatbig], which was published in spring 2025—after the release of all of the models we test. (The newest model is from the Qwen 2.5 series, which was released in late September 2024.) This book is not in any of the models’ training data. We run probabilistic extraction on (a) Llama 3.1 70B, (b) Llama 3.1 8B, (c) Llama 2 13B, (d) Phi 4, and (e) Qwen 2.5 72B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. We are unable to extract any unique (i.e., non-copyright notice) sequences using any model or prefix length.